OpenStudy (anonymous):

prove that cos 20 * cos 40 * cos 80 = 1/8

OpenStudy (anonymous):

Which of the following best describes the graph of f(x) = x ^2 - 7x + 10? Minimum at (1.5, -12.25) with intercepts at (5, 0) and (-2, 0) Minimum at (-1.5, -12.25) with intercepts at (-5, 0) and (2, 0) Minimum at (-3.5, -2.25) with intercepts at (-5, 0) and (-2, 0) Minimum at (3.5, -2.25) with intercepts at (5, 0) and (2, 0)

OpenStudy (anonymous):

poop

OpenStudy (anonymous):

i need help here not poop

OpenStudy (anonymous):

L.H.S = cos 20 * cos 40 * cos 80 = 2*sin(20)*cos20*cos40*cos80 / (2sin20) = sin(40)*cos(40)*cos(80)/(2sin20) = sin(80)*cos(80)/(4sin20) =sin160 / ( 8 sin20) = sin(180-20)/(8sin20) =sin20/(8sin20) =1/8 = R.H.S